# Multi-Group Proportional Representation

### v1

@@ Abstract
Image search and retrieval tasks can perpetuate harmful stereotypes, erase cultural identities, and amplify social disparities. Current approaches to mitigate these representational harms balance the number of retrieved items across population groups defined by a small number of (often binary) attributes. However, most existing methods overlook intersectional groups determined by combinations of group attributes, such as gender, race, and ethnicity. We introduce Multi-Group Proportional Representation (MPR), a novel metric that measures representation across intersectional groups. We develop practical methods for estimating MPR, provide theoretical guarantees, and propose optimization algorithms to ensure MPR in retrieval. We demonstrate that existing methods optimizing for equal and proportional representation metrics may fail to promote MPR. Crucially, our work shows that optimizing MPR yields more proportional representation across multiple intersectional groups specified by a rich function class, often with minimal compromise in retrieval accuracy.
@@ 1 Introduction
We focus on representation in retrieval tasks. Representational harms in retrieval and ranking tasks arise when retrieved items do not accurately reflect the true diversity and proportions present in reality, perpetuating harmful stereotypes and biases [13]. For instance, Kay et al. [2] demonstrated that, in 2015, only $10\%$ of the top 100 results for CEO in Google Image Search were women, despite $28\%$ of CEOs in the US being women, and recent studies [14, 15] have demonstrated that this bias is still present in modern image search engines. Later, Otterbacher et al. [16] showed that the search engine Bing usually produced twice as many men as compared to women when queried with the word “person”. Finally, [17] systematically studied gender biases in image representations by auditing the four most important image search engines: Google, Bing, Baidu, and Yandex. In particular, among other findings, this study showed that when a qualifier such as “intelligent” is added to the query “person”, the engines still exhibit a huge gender discrepancy; see also [1, Chapter 7]. When retrieval is based on a search over vector embeddings of images and text, biases in embedding models can propagate to retrieved results, leading to disparate representation of people by demographic groups such as gender or race [18, 19, 20], as well as propagating spurious correlations from the dataset more broadly [21] or distorting results due to stereotypes present in the representation [22]. Though correlations between unrelated semantic concepts do not necessarily constitute representational harms, a wealth of research charts undesirable bias in embedding models, including publicly available joint vision-language foundational embedding models such as CLIP [23] used for retrieval. Bias in embeddings can lead to bias in retrieval [24, 25, 26]. In fact, recent work by Srinivasan et al. [27] argues that many image retrieval systems lack “social diversity” as perceived by humans.
@@
Several interventions aim to control gross statistical deviations in group representations in retrieval. A common approach is ensuring equal representation of pre-defined population groups in retrieved items [28, 29]. An alternative goal is proportional representation (PR) [30, 20], where the representation of different groups is approximately equal to reference population statistics. Various interventions have been proposed to achieve equal or proportional representation, including optimizing for diversity in addition to the similarity in retrieval [31, 27], directly selecting the objects retrieved based on balancing known or predicted sensitive attributes [29] and, for vector databases, modifying embeddings directly to remove information about group-defining attributes [18, 20].
@@
Achieving proportional representation across multiple intersectional groups is challenging, as existing fairness interventions typically consider only a small number of pre-defined groups. Ensuring representation across individual groups (e.g., given by gender or race) does not guarantee representation across intersectional groups (e.g., given by gender and race), as demonstrated in Tables 1 and 3. The study of intersectionality and its consequences is rooted in work in law and the social sciences [32, 33, 34] and has been a well-discussed problem for several decades. In machine learning systems, lack of intersectional representation can contribute to the invisibility of historically marginalized groups determined by multiple axes of identity [35], or their mistreatment through “fairness gerrymandering,” where interventions on fairness for specific groups may harm fairness on intersectional subgroups [36]. However, as the number of group-denoting attributes increases, the number of intersectional groups grows exponentially, quickly surpassing the number of retrieved items and limiting the achievable proportional representation.
@@
We propose a metric called Multi-group Proportional Representation (MPR) to quantify the representation of intersectional groups in retrieval tasks. MPR measures the worst-case deviation between the average values of a collection of representation statistics computed over retrieved items relative to a reference population whose representation we aim to match. The set of representation statistics is given by a function class $\mathcal{C}$, where each function $c\in\mathcal{C}$ maps a retrieved item $x$ to a Real number. For instance, $c(x)\in\{-1,1\}$ may denote binary group membership. In theory, $\mathcal{C}$ can be given by bounded complexity function classes, such as linear functions or shallow decision trees. In practice, $\mathcal{C}$ are functions defined over item attributes, such as vector embeddings or group-denoting labels, enabling MPR to measure proportional representation across complex, intersectional subgroups. Compared to naively counting the number of retrieved items across a (potentially exponential) number of pre-defined groups, MPR offers a more flexible, scalable, and theoretically grounded metric for multi-group representation in retrieval.
@@
In addition to introducing and carefully motivating the MPR metric (Section 2), we make three contributions. First, we show how to compute MPR for several function classes in Section 3. Computing MPR relies on a curated dataset that represents the desired population whose proportional representation we aim to reflect in retrieved items. We derive sample complexity bounds for the size of this curated dataset based on the complexity of the function class $\mathcal{C}$. We also demonstrate that MPR can be computed in closed form for certain function classes or with a regression oracle. Second, we propose MAPR, an algorithm that retrieves items from a vector database to maximize their average similarity to a query embedding while satisfying an MPR constraint relative to the curated dataset (Section 4). MAPR achieves this by iteratively calling an oracle that computes MPR violations, yielding both relevant and representative retrieved items. Third, we evaluate MAPR on retrieval tasks using datasets such as CelebA [37], the Occupations dataset [31], Fairface [38], and UTKFace [39]. Our results (Section 5) show that MAPR Pareto-dominates competing approaches in balancing retrieval similarity and MPR.
@@
Mitigating Bias in Embedding Space. Various approaches have been proposed to produce vector embeddings of text and images that target diversity or mitigate bias. These include modifying loss functions to encourage group fairness [40], adversarial training [41, 42], disentangling representations [43], and in-processing fair sampling methods for gender imbalance [18]. Many foundational embeddings models are not trained with such approaches, so post-processing methods for fairness have also been developed for multimodal models like CLIP [23], such as CLIP-clip [18], CLIP-debias [20], and FairCLIP [19]. The recent work PATHS [27], built upon the pretrained image-text model CoCa, leverages text-guided projections to extract an embedding that captures complex notions of diversity in people, enabling effective diverse retrieval methods. In our experiments, we report the MPR achieved by CLIP-clip and CLIP-debias in retrieval tasks. PATHS [27] is the recent work most similar to ours in that it also aims to promote intersectional representation in retrieval. Code for implementing and reproducing [27] was not available at the time of submission. We approximate its retrieval algorithm with vanilla CLIP embeddings. Unlike the aforementioned methods, MAPR is not based on modifying embeddings.
@@
Retrieval Algorithms with Diversity or Fairness. Instead of modifying embeddings, several methods aim to promote diversity, given possibly biased embeddings. For a fixed query, the baseline method for retrieval is $k$-Nearest-Neighbors ($k$-NN), where the $k$ most similar items are retrieved [44]. Given an additional diversity constraint, a popular approach is Maximal Marginal Relevance (MMR) [45], where items are greedily retrieved in order to maximize a linear combination of similarity and diversity metrics. Celis et al. [30], Celis and Keswani [31] focus on diversity and fairness in image retrieval, using a reweighing method that combines similarity and MMR to select diverse yet relevant images. Alternatively, the Post-Hoc Bias Mitigation [29] method aims to achieve equal representation over a pre-defined attribute (such as gender) through calls to an oracle classifier and selecting a balanced group of images. We extend this prior work in two key ways: (i) we aim to achieve proportional representation across intersectional groups, going beyond the representation of two groups defined by binary attributes; (ii) MAPR explicitly balances MPR and retrieval similarity via an optimization that controls for both, offering a principled alternative to greedy MMR-based methods.
@@
Diversity Metrics. A traditional measure of diversity for a set is the pairwise similarity of the retrieved items [46]. However, the general visual (dis)similarity metric is insufficient in optimizing for people-related diversity such as skin tone or gender [30]. When the attribute (e.g., race) along which diversity is optimized is known, fairness definition-derived metrics such as proportional representation are used ([47], see also next section). For more complex notions of diversity that aim to capture intersectional or subtle sociocultural identities (e.g., cultural backgrounds, lifestyles, nationalities, religions), human annotators are employed to determine if the selected set of images is more diverse than the baseline (retrieval based on similarity to query only) [27]. MPR complements existing metrics by providing a rigorous approach to measuring representation across complex, intersectional groups defined by a rich class of functions. However, as with any fairness metric, MPR has limitations (see Section 6) and may not always align with human perceptions of representation. We underscore the importance of involving stakeholders in defining representation goals and auditing retrieval systems.
@@
Multi-Group Fairness. Our work is directly informed by the burgeoning literature on multi-group fairness in classification, particularly the work of Hébert-Johnson et al. [48], Kim et al. [49, 50] who proposed rigorous frameworks of multi-group fairness auditing and post-processing to ensure fair predictions across identifiable subgroups. Recent efforts in this field include [51, 52, 53], which analyze sample complexity aspects of measuring and reporting multi-group representation, and [54, 55] who develop stronger variants of multi-group fairness notions. Unlike prior work, we focus on multi-group representation in retrieval rather than multi-group fairness in classification and regression. Lastly, proportional representation strategies arose in the political sciences literature [56, 57, 58, 59, 60] – we address this body of work in Appendix C.
@@ 2 A Multi-Group Proportional Representation Metric
With a group-denoting function $c(\boldsymbol{x}^{r})$ in hand, we can measure the representation of each group in a set of retrieved items $\mathcal{R}(\mathbf{q})$. One popular constraint for representation is equal representation [28, 2], which aims to ensure that the number of retrieved items in each group is approximately the same, i.e., $\frac{1}{k}\sum_{\boldsymbol{x}^{r}\in\mathcal{R}(\mathbf{q})}c(\boldsymbol{x}^{r})\approx 0$. An alternative metric is proportional representation [61, 20], which quantifies the deviation of group membership from a reference distribution $Q$. Methods that promote proportional representation aim to ensure $\frac{1}{k}\sum_{\boldsymbol{x}^{r}\in\mathcal{R}(\mathbf{q})}c(\boldsymbol{x}^{r})\approx\mathbb{E}_{Q}\left[c(X)\right]$.
Here, the measure $Q$ captures the distribution of a reference population whose representation statistics we aim to match. For example, if $Q$ were the distribution of individuals in the United States, $\mathbb{E}_{Q}\left[c(X)\right]$ could measure the proportion of men vs. women in the US and be approximated using Census data.
@@
The MPR metric quantifies the “representativeness” of a rich class of statistics, denoted by functions in $\mathcal{C}$, within a set of retrieved items $\mathcal{R}$. This generalization is crucial for capturing intersectional groups. For example, $\mathcal{C}$ could contain functions that map items to demographic groups based on race, gender, age, or combinations thereof. Alternatively, when labels in $\mathcal{G}$ contain such group-denoting attributes, $\mathcal{C}$ can represent decision trees of a given depth over features (e.g., all combinations of pairs of race, gender, and age). The MPR metric compares the empirical average of each $c\in\mathcal{C}$ over the retrieved set $\mathcal{R}$ to its expectation under the reference distribution $Q$. By defining $Q$ appropriately, we can flexibly specify different statistical representation goals, such as equal representation ($\mathbb{E}_{Q}[c(X)]=1/2$ for binary $c$) or proportional representation w.r.t. a target population. However, as $\mathcal{C}$ grows in complexity, pre-specifying $\mathbb{E}_{Q}[c(X)]$ for all $c\in\mathcal{C}$ becomes infeasible. Instead, MPR can be estimated using a carefully curated dataset, which we define in the next subsection.
@@
The exact nature of the curated dataset $\mathcal{D}_{C}$ is context-dependent and may vary based on the specific application and desired representation goals. We can also condition the curated dataset on a given query. Specifically, for a query $\mathbf{q}$, we can retrieve relevant samples from both the curated dataset $\mathcal{D}_{C}$ and the retrieval dataset $\mathcal{D}_{R}$. The samples retrieved from $\mathcal{D}_{C}$ can then serve as a “conditional” curated dataset, denoted as $\mathcal{D}_{C|\mathbf{q}}$, which captures the desired representation target specific to the query $\mathbf{q}$. This approach allows for a more granular and context-aware proportional representation.
In the next two sections, we introduce theoretical guarantees and algorithms that are agnostic to the specific choice of $\mathcal{D}_{C}$. In other words, our proposed methods for measuring and promoting MPR in retrieval are general and can be applied regardless of how the curated dataset is constructed. In our numerical experiments (Section 5), we use FairFace [38] as the curated dataset.

### v2

@@ Abstract
Image search and retrieval tasks can perpetuate harmful stereotypes, erase cultural identities, and amplify social disparities. Current approaches to mitigate these representational harms balance the number of retrieved items across population groups defined by a small number of (often binary) attributes. However, most existing methods overlook intersectional groups determined by combinations of group attributes, such as gender, race, and ethnicity. We introduce Multi-Group Proportional Representation (MPR), a novel metric that measures representation across intersectional groups. We develop practical methods for estimating MPR, provide theoretical guarantees, and propose optimization algorithms to ensure MPR in retrieval. We demonstrate that existing methods optimizing for equal and proportional representation metrics may fail to promote MPR. Crucially, our work shows that optimizing MPR yields more proportional representation across multiple intersectional groups specified by a rich function class, often with minimal compromise in retrieval accuracy. Code is provided at https://github.com/alex-oesterling/multigroup-proportional-representation.
footnotetext: Equal contribution, correspondence to aoesterling@g.harvard.edu
@@ 1 Introduction
We focus on representation in retrieval tasks. Representational harms in retrieval and ranking tasks arise when retrieved items do not accurately reflect the true diversity and proportions present in reality, perpetuating harmful stereotypes and biases [13]. For instance, Kay et al. [2] demonstrated that, in 2015, only $10\%$ of the top 100 results for CEO in Google Image Search were women, despite $28\%$ of CEOs in the US being women, and recent studies [14, 15] have demonstrated that this bias is still present in modern image search engines. Later, Otterbacher et al. [16] showed that the search engine Bing usually produced twice as many men as compared to women when queried with the word “person”. Finally, [17] systematically studied gender biases in image representations by auditing the four most important image search engines: Google, Bing, Baidu, and Yandex. In particular, among other findings, this study showed that when a qualifier such as “intelligent” is added to the query “person”, the engines still exhibit a significant gender discrepancy; see also [1, Chapter 7]. When retrieval is based on a search over vector embeddings of images and text, biases in embedding models can propagate to retrieved results, leading to disparate representation of people by demographic groups such as gender or race [18, 19, 20], as well as propagating spurious correlations from the dataset more broadly [21] or distorting results due to stereotypes present in the representation [22]. Though correlations between unrelated semantic concepts do not necessarily constitute representational harms, a wealth of research charts undesirable bias in embedding models, including publicly available vision-language embedding models such as CLIP [23] used for retrieval. Bias in embeddings can lead to bias in retrieval [24, 25, 26]. In fact, recent work by Srinivasan et al. [27] argues that many image retrieval systems lack “social diversity” as perceived by humans.
@@
Several interventions aim to control gross statistical deviations in group representations in retrieval. A common approach is ensuring equal representation of pre-defined population groups in retrieved items [28, 29]. An alternative goal is proportional representation (PR) [30, 20], where the target representation of different groups is proportional to some reference population or statistic. Various interventions have been proposed to achieve equal or proportional representation, including optimizing for diversity in addition to the similarity in retrieval [31, 27], directly selecting the objects retrieved based on balancing known or predicted sensitive attributes [29] and, for vector databases, modifying embeddings directly to remove information about group-defining attributes [18, 20].
@@
Achieving proportional representation across multiple intersectional groups is challenging, as existing fairness interventions typically consider only a small number of pre-defined groups. Ensuring representation across individual groups (e.g., given by gender or race) does not guarantee representation across intersectional groups (e.g., given by gender and race), as demonstrated in Table 1 (and Table 5 in the Appendix). The study of intersectionality and its consequences is rooted in work in law and the social sciences [32, 33, 34] and has been a well-discussed problem for several decades. In machine learning systems, lack of intersectional representation can contribute to the invisibility of historically marginalized groups determined by multiple axes of identity [35], or their mistreatment through “fairness gerrymandering,” where interventions on fairness for specific groups may harm fairness on intersectional subgroups [36]. However, as the number of group-denoting attributes increases, the number of intersectional groups grows exponentially, quickly surpassing the number of retrieved items and limiting the achievable proportional representation. In fact, achieving optimal proportional representation across multiple attributes has been shown to be NP-hard [37].
@@
We propose a metric called Multi-group Proportional Representation (MPR) to quantify the representation of intersectional groups in retrieval tasks. MPR measures the worst-case deviation between the average values of a collection of representation statistics computed over retrieved items relative to a reference population whose representation we aim to match. The set of representation statistics is given by a function class $\mathcal{C}$, where each function $c\in\mathcal{C}$ maps a retrieved item $x$ to a real number. For instance, $c(x)\in\{-1,1\}$ may denote binary group membership for an input $x$. In theory, $\mathcal{C}$ can be given by bounded complexity function classes, such as linear functions or shallow decision trees. In practice, $\mathcal{C}$ are functions defined over item attributes, such as vector embeddings or group-denoting labels, enabling MPR to measure proportional representation across complex, intersectional subgroups. Compared to naively counting the number of retrieved items across a (potentially exponential) number of pre-defined groups, MPR offers a more flexible, scalable, and theoretically grounded metric for multi-group representation in retrieval.
@@
In addition to introducing and carefully motivating the MPR metric, cf. Section 2, we make three contributions. First, we show how to compute MPR for several function classes in Section 3. Computing MPR relies on a curated dataset that represents the desired population whose proportional representation we aim to reflect in retrieved items. We derive sample complexity bounds for the size of this curated dataset based on the complexity of the function class $\mathcal{C}$. We also prove that MPR can be computed in closed form for certain function classes or with a regression oracle. Second, we propose MOPR, an algorithm that retrieves items from a vector database to maximize their average similarity to a query embedding while satisfying an MPR constraint relative to the curated dataset (Section 4). MOPR achieves this by iteratively calling an oracle that computes MPR violations, yielding both relevant and representative retrieved items. Third, we evaluate MOPR on retrieval tasks using datasets such as CelebA [38], the Occupations dataset [31], Fairface [39], and UTKFace [40]. In section 5, we show that MOPR Pareto-dominates competing approaches in balancing retrieval similarity and MPR.
@@
Mitigating Bias in Embedding Space. Various approaches have been proposed to produce vector embeddings of text and images that target diversity or mitigate bias. These include modifying loss functions to encourage group fairness [41], adversarial training [42, 43], disentangling representations [44], and in-processing fair sampling methods for gender imbalance [18]. Many popular embedding models are not trained with such approaches, so post-processing methods for fairness have also been developed for multimodal models such as CLIP [23], including CLIP-clip [18], CLIP-debias [20], and FairCLIP [19]. Recent work by Srinivasan et al. [27] builds upon the pretrained image-text model CoCa and leverages text-guided projections to extract an embedding, called PATHS, that captures complex notions of diversity in people, and then conduct diverse retrieval with a greedy algorithm. However, code for implementing and reproducing PATHS embeddings was not available at the time of publication. In our experiments, we report the MPR achieved by CLIP-clip and CLIP-debias in retrieval tasks and approximate the greedy retrieval algorithm proposed in [27] with vanilla CLIP embeddings. Unlike the aforementioned methods, MOPR is not based on modifying embeddings.
@@
Retrieval Algorithms with Diversity or Fairness. Instead of modifying embeddings, several methods aim to promote diversity, given possibly biased embeddings. Given an additional diversity constraint, a popular approach is Maximal Marginal Relevance (MMR) [45], where items are greedily retrieved in order to maximize a linear combination of similarity and diversity metrics. Celis et al. [30] and Celis and Keswani [31] focus on diversity and fairness in image retrieval, using a reweighing method that combines similarity and MMR to select diverse yet relevant images. Alternatively, the Post-Hoc Bias Mitigation [29] method aims to achieve equal representation over a pre-defined attribute (such as gender) through calls to an oracle classifier and selecting a balanced group of images. We extend this prior work in two key ways: (i) we aim to achieve proportional representation across intersectional groups, going beyond the representation of two groups defined by binary attributes; (ii) MOPR explicitly balances MPR and retrieval similarity via an optimization that controls for both, offering an efficient alternative to greedy MMR-based methods.
@@
Diversity Metrics. A traditional measure of diversity for a set is the pairwise similarity of the retrieved items [46]. However, the general visual (dis)similarity metric is insufficient in optimizing for people-related diversity such as skin tone or gender [30]. When the attribute (e.g., race) along which diversity is optimized is known, or, in other words, we have access to a ground truth attribute label, one can use fairness definition-derived metrics such as proportional representation (see next section and [47]). For more complex notions of diversity that aim to capture intersectional or subtle sociocultural identities (e.g., cultural backgrounds, lifestyles, nationalities, religions), human annotators can be employed to determine if the selected set of images is more diverse than the baseline (retrieval based on similarity to query only) [27]. MPR complements existing metrics by providing a rigorous approach to measuring representation across complex, intersectional groups defined by a rich class of functions. However, as with any fairness metric, MPR has limitations – see Section 6 – and may not always align with human perceptions of representation. In particular, we underscore the importance of involving stakeholders in defining representation goals and auditing retrieval systems.
@@
Multi-Group Fairness. Our work is directly informed by the burgeoning literature on multi-group fairness in classification, particularly the work of Hébert-Johnson et al. [48], Kim et al. [49, 50] who proposed rigorous frameworks of multi-group fairness auditing and post-processing to ensure fair predictions across identifiable subgroups. Recent efforts in this field include [51, 52, 53], which analyze sample complexity aspects of measuring and reporting multi-group representation, and [54, 55] who develop stronger variants of multi-group fairness notions. Unlike prior work, we focus on multi-group representation in retrieval rather than multi-group fairness in classification and regression.
Multi-Attribute Proportional Representation. A recognized challenge in computational social choice research is building committees, groups, or sets of items that ensure proportional representation across several attributes [56, 57, 58, 59]. The work most closely related to ours is by Lang and Skowron [37], which introduced the problem of multi-attribute proportional representation, where a set of items must be selected to reflect desired distributions over multiple attributes simultaneously. Their work established fundamental computational complexity results, showing that finding optimal proportional representation is NP-hard, and developed approximation algorithms and integer linear programs for promoting proportional representation across attributes. While our work shares similar goals, our MPR metric takes a different approach by measuring the worst-case deviation over computationally-identifiable groups defined by a function class $\mathcal{C}$. This formulation connects naturally with statistical learning theory through MMD [60] and sample complexity bounds (Propositions 1 and 2). Though we use the term multi-group rather than multi-attribute to emphasize connections with the burgeoning literature on multi-group fairness [48, 49], both terms describe related problems of ensuring representation across multiple dimensions. We also highlight that work on multi-attribute representation in computational social choice (including [37]) strives for proportional representation marginally on each attribute, while our primary goal with this work is to achieve intersectional representation. Finally, while [37] similarly seeks to match retrieved candidates to a target distribution, we focus on database retrieval rather than social choice applications. The rich history of proportional representation in social choice theory and political sciences provides, however, an important context for our work. We briefly discuss these connections in Appendix E. There, we elaborate on how our approach relates to and differs in terms of representation formulation from [37].
@@ 2 A Multi-Group Proportional Representation Metric
With a group-denoting function $c(\boldsymbol{x}^{r})$ in hand, we can measure the representation of each group in a set of retrieved items $\mathcal{R}(\mathbf{q})$. One popular constraint for representation is equal representation [28, 2], which aims to ensure that the number of retrieved items in each group is approximately the same, i.e., $\frac{1}{k}\sum_{\boldsymbol{x}^{r}\in\mathcal{R}(\mathbf{q})}c(\boldsymbol{x}^{r})\approx 0$. An alternative metric is proportional representation [61, 20], which quantifies the deviation of group membership from a reference distribution $Q$. Methods that promote proportional representation aim to ensure $\frac{1}{k}\sum_{\boldsymbol{x}^{r}\in\mathcal{R}(\mathbf{q})}c(\boldsymbol{x}^{r})\approx\mathbb{E}_{Q}\left[c(X)\right]$. Here, the measure $Q$ captures the distribution of a reference population whose representation statistics we aim to match. For example, if $Q$ were the distribution of individuals in the US, $\mathbb{E}_{Q}\left[c(X)\right]$ could measure the proportion of men vs. women in the US and be approximated using Census data.
@@
The MPR metric quantifies the “representativeness” of a rich class of statistics, denoted by functions in $\mathcal{C}$, within a set of retrieved items $\mathcal{R}$. This generalization is crucial for capturing intersectional groups. For example, $\mathcal{C}$ could contain functions that map items to demographic groups based on race, gender, age, or combinations thereof. Alternatively, when labels in $\mathcal{G}$ contain such group-denoting attributes, $\mathcal{C}$ can represent decision trees of a given depth over features (e.g., all combinations of pairs of race, gender, and age). The MPR metric compares the empirical average of each $c\in\mathcal{C}$ over the retrieved set $\mathcal{R}$ to its expectation under the reference distribution $Q$. By defining $Q$ appropriately, we can flexibly specify different statistical representation goals, such as equal representation ($\mathbb{E}_{Q}[c(X)]=0$ for binary $c$) or proportional representation w.r.t. a target population. Measuring representation requires defining what constitutes “fair and proportional representation.” While equal representation may suffice for binary groups, proportional representation of more complex intersectional groups requires care. The choice of representation reference statistics (given by the distribution Q in the definition of MPR) should be application-dependent, context-aware, and culturally sensitive.
@@
Constructing a proper $\mathcal{D}_{C}$ is critical to properly measuring bias and preventing downstream harms, and the nature of $\mathcal{D}_{C}$ is context-dependent and may vary based on the specific application and desired representation goals. We reiterate that if the curated dataset is biased, then these biases will be propagated to downstream usages of MPR. For these reasons, we recommend that curated datasets used for MPR measurements be developed and verified through participatory design approaches [66, 67] in collaboration with diverse stakeholders. By involving stakeholders in defining representation goals, we can help ensure that the proportional representation in information retrieval systems measured by MPR aligns with the values and needs of the user base these systems serve.
Finally, we note that we can also condition the curated dataset on a given query. Specifically, for a query $\mathbf{q}$, we can retrieve relevant samples from both the curated dataset $\mathcal{D}_{C}$ and the retrieval dataset $\mathcal{D}_{R}$. The samples retrieved from $\mathcal{D}_{C}$ can then serve as a “conditional” curated dataset, denoted as $\mathcal{D}_{C|\mathbf{q}}$, which captures the desired representation target specific to the query $\mathbf{q}$. This approach allows for a more granular and context-aware proportional representation.
In the next two sections, we introduce theoretical guarantees and algorithms that are agnostic to the specific choice of $\mathcal{D}_{C}$. In other words, our proposed methods for measuring and promoting MPR in retrieval are general and can be applied regardless of how the curated dataset is constructed. In our numerical experiments, presented in Section 5, we use FairFace [39] as the curated dataset.